9. A products demand over (/+ 1) periods follows a normal distribution with mean of 80 and standard deviation of 20. The order-up-to level is 100. What is the in-stock probability? 01 O 0.8413 。0.5987 O 0.3413
Answer #1

Ans: 0.8413

The order level is 100.So the order upto the level of 100 can be fulfilled at one go.

Now the demand follows the normal probability distribution curve.

Here Z= (X- mean) /(Standard deviation) = (100-80)/20 = 1.00

Now looking for the probability from Z table we get value = 0.8413

2 000 001 0.02 0.03 0.04 005 0.06 007 0.08 0.09 00 5000 5040 5080 5120 5160 51995239 5279 5319 5359 0.1 5338 5438 5478 5517 5